### v1

@@ 1 Introduction
If the maximum frequency $f$ (the maximum number of sets that any element belongs to) is bounded, we can drop the assumption that $k=\Omega(m)$, and parameterize the number of rounds based on $f$. In particular, we can obtain a $1-1/e-\epsilon$ approximation in $O(f^{2}/\epsilon^{4}\cdot\log m\cdot(\log(1/\epsilon)+\log m))$ rounds.
@@ 2.1 The main algorithm
$\sum_{j=1}^{m}y^{\prime}_{i}$ $=\sum_{j=1}^{m}\left(1-\frac{z_{i}}{1+\epsilon}\right)=m-\frac{m-k}{1+\epsilon}\leq m-(1-2\epsilon)(m-k)<k+2\epsilon m,$ (2)
@@
We set $\epsilon=\eta^{2}/f$ and proceed to obtain a solution that covers at least $(1-\eta^{2}/f)(1-\eta)\textup{OPT}=(1-O(\eta))\textup{OPT}$ elements using at most $k+O(\epsilon m)=k+O(\eta^{2}/f\cdot kf/\eta)=k+O(\eta k)$ sets as in the discussion above. Appealing to Lemma 3, we can find $k$ sets that cover at least $(1-O(\eta))\textup{OPT}$ elements. The total number of rounds is $O(1/\eta^{4}\cdot\log m\cdot(\log(1/\epsilon)+\log m))$.
@@ 2.2 Approximate the LP’s solution via multiplicative weights
3       Run the oracle in Lemma 4 with $\textbf{w}^{(t-1)}$ to check if there exists a feasible solution. If the answer is Infeasible, stop the algorithm. If the answer is Feasible, let $\textbf{x}^{(t)}$ and $\textbf{z}^{(t)}$ be the output of the oracle that are now stored in all machines. .
@@
8       For each $i\in[n]$, each machine computes $w_{i}^{(t)}=2^{-\epsilon\cdot\sum_{d=1}^{t}E_{i}^{(d)}/f_{i}}=2^{-\epsilon\cdot\sum_{d=1}^{t}\textup{error}_{i}^{(d)}}=2^{-\epsilon\cdot\textup{error}_{i}^{(t)}}w_{i}^{(t-1)}$.
@@
10After $T$ iterations, output $\textbf{x}=\frac{1}{T}\sum_{t=1}^{T}\textbf{x}^{(t)}$ and $\textbf{z}=\frac{1}{T}\sum_{t=1}^{T}\textbf{z}^{(t)}$.

### v2

@@ 1 Introduction
If the maximum frequency $f$ (the maximum number of sets that any element belongs to) is bounded, we can drop the assumption that $k=\Omega(m)$, and parameterize the number of rounds based on $f$. In particular, we can obtain a $1-1/e-\epsilon$ approximation in $O(f^{3}/\epsilon^{6}\cdot\log^{2}(\frac{kf}{\epsilon}))$ rounds.
@@ 2.1 The main algorithm
$\sum_{j=1}^{m}y^{\prime}_{i}$ $=\sum_{j=1}^{m}\left(1-\frac{z_{j}}{1+\epsilon}\right)=m-\frac{m-k}{1+\epsilon}\leq m-(1-2\epsilon)(m-k)<k+2\epsilon m,$ (2)
@@
We set $\epsilon=\eta^{2}/f$ and proceed to obtain a solution that covers at least $(1-\eta^{2}/f)(1-\eta)\textup{OPT}=(1-O(\eta))\textup{OPT}$ elements using at most $k+O(\epsilon m)=k+O(\eta^{2}/f\cdot kf/\eta)=k+O(\eta k)$ sets as in the discussion above. Appealing to Lemma 3, we can find $k$ sets that cover at least $(1-O(\eta))\textup{OPT}$ elements. The total number of rounds is $O(f^{3}/\eta^{6}\cdot\log\frac{kf}{\eta}\cdot(\log\frac{1}{\eta}+\log\frac{kf}{\eta}))=O(f^{3}/\eta^{6}\cdot\log^{2}\frac{kf}{\eta})$.
@@ 2.2 Approximate the LP’s solution via multiplicative weights
3       Run the oracle in Lemma 4 with $\textbf{w}^{(t-1)}$ to check if there exists a feasible solution. If the answer is Infeasible, stop the algorithm. If the answer is Feasible, let $\textbf{x}^{(t)}$ and $\textbf{z}^{(t)}$ be the output of the oracle that are now stored in all machines .
@@
8       For each $i\in[n]$, each machine computes
$w_{i}^{(t)}=2^{-\epsilon\cdot\sum_{d=1}^{t}E_{i}^{(d)}/f_{i}}=2^{-\epsilon\cdot\sum_{d=1}^{t}\textup{error}_{i}^{(d)}}=2^{-\epsilon\cdot\textup{error}_{i}^{(t)}}w_{i}^{(t-1)}.$
@@ 8       For each $i\in[n]$, each machine computes
10After $T$ iterations, output
$\textbf{x}=\frac{1}{T}\sum_{t=1}^{T}\textbf{x}^{(t)}\text{, and }\textbf{z}=\frac{1}{T}\sum_{t=1}^{T}\textbf{z}^{(t)}.$